Circular No. 7304

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999et IN NGC 1643 E. Cappellaro, Osservatorio Astronomico di Padova, reports his discov-

ery, on an R CCD frame obtained with the Danish telescope (+ DFOSC) at the European Southern Observatory, La Silla, on Nov. 4.3 UT, of a supernova (R = 17.6) located at α = 4^{h}43^{m}44^{s}.0, δ = −5^{o}19′09′′ (equinox 2000.0), which is 5′′.8 west and 9′′.8 south of the nucleus of NGC 1643. A spectrum obtained with same instrument (range 330–900 nm, resolution 1 nm) shows that the object is a typical type-II supernova, 1–2 months after maximum. The expansion velocity measured from the well-developed P-Cyg absorption of Hα is ∼ 6000 km/s.

SUPERNOVA 1999eu IN NGC 1097 S. Nakano, Sumoto, Japan, reports the discovery by Masakatsu Aoki,

Toyama, of a supernova (mag 17.3) on seventeen unfiltered CCD frames taken on Nov. 5 with a 0.40-m f/6 reflector, with confirming frames taken on Nov. 6. Previous frames taken by Aoki in 1996, 1997, and 1998 show nothing at this location (limiting mag 18–19). SN 1999eu is located at α = 2^{h}46^{m}20^{s}.79, δ = −30^{o}19′06′′.1 (equinox 2000.0), which is 23′′ east and 157′′ south of the center of NGC 1097.

P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics; and B. Schmidt, Mount Stromlo and Siding Springs Observatory, write: “A spectrum was obtained with the Fred L. Whipple Observatory 1.5-m Tillinghast telescope by P. Berlind on Nov. 8.3 UT, which shows that SN 1999eu is a peculiar type-II supernova. A large number of narrow P-Cyg features are distributed through the spectrum, and Hα has strong absorption bands superimposed on its profile. Na I 589.3nm shows a well-developed P-Cyg profile, as does Ba II 614.2-nm. The expansion velocity measured from the Ba II line is only 1500 km/s. The spectrum is nearly identical to that of SN 1997D (Turatto et al. 1998, Ap.J. Let. 498, 129), a subluminous type-II event that is suspected of ejecting an extremely small amount of ^{56}Ni.”

SUPERNOVAE 1999dq, 1999eh, 1999ej, 1999el R magnitudes by K. Sárneczky, L. Kiss, and G. Szabó, at Piszkéstető,

Konkoly Observatory: SN 1999dq in NGC 976, Oct. 21.989 UT, 16.2 ± 0.3; SN 1999eh in NGC 2770, Oct. 26.069, 18.3 ± 0.3; SN 1999ej in NGC 495, Oct. 27.742, 16.1 ± 0.1; SN 1999el in NGC 6951, Oct. 27.709, 15.0 ± 0.2.

1999 November 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green